### v1

# Two-dimensional fractional Brownian motion: Analysis in time and frequency domains

Abstract
This article introduces a novel construction of the two-dimensional fractional Brownian motion (2D fBm) with dependent components. Unlike similar models discussed in the literature, our approach uniquely accommodates the full range of model parameters and explicitly incorporates cross-dependencies and anisotropic scaling through a matrix-valued Hurst operator. We thoroughly analyze the theoretical properties of the proposed causal and well-balanced 2D fBm versions, deriving their auto- and cross-covariance structures in both time and frequency domains. In particular, we present the power spectral density of these processes and their increments. Our analytical findings are validated with numerical simulations. This work provides a comprehensive framework for modeling anomalous diffusion phenomena in multidimensional systems where component interdependencies are crucial.
2D fBm, vector fractional Brownian motion, power spectral density, cross spectral density, multivariate self-similar processes, autocorrelation function, cross-correlation function
I Introduction
The study of anomalous diffusion phenomena has emerged as a critical area of research in diverse scientific fields, from cellular biophysics to financial mathematics [34, 24, 31, 47, 43, 37]. While standard Brownian motion adequately describes normal diffusion processes characterized by a mean-square displacement that increases linearly with time [14], multiple complex systems exhibit a non-linear scaling that demands a more sophisticated mathematical framework [34]. Fractional Brownian motion (fBm), formalized by Mandelbrot and van Ness extends the classical Brownian motion model by introducing long-range dependence through the Hurst parameter $H$ [30]. However, when considering multidimensional systems, the scalar parameter $H$ can be insufficient to capture the complex cross-dependencies and anisotropic scaling behaviors observed in empirical data. Multidimensional fBm or vector fBm [33, 20, 28, 1], thus, represents an advancement in the modeling of such anomalous diffusion processes in multidimensional settings.
A common approach to modeling multidimensional fBm involves treating each spatial component as an independent random walk. This strategy has been employed in a variety of contexts, including the analysis of diffusion in living cells and complex fluids, where isotropic behavior is often assumed over the observation timescales [48, 25, 50, 36]. For instance, models with independent fBm components have been used to characterize anomalous transport of membrane proteins [48], analyze motion changes in single-particle trajectories [36], and assess spectral properties of anomalous diffusion processes [50, 36]. While this approach is well justified in systems where anisotropies are either negligible or not detectable within the experimental resolution, it becomes insufficient in environments where directional dependence or cross-correlations play a significant role. In such cases, assuming independence between components can obscure critical features of the dynamics and lead to misinterpretations. This motivates the development of models that explicitly incorporate cross-dependencies and anisotropic scaling.
While one expects many biophysical systems to be isotropic [8, 59, 44], this is not the case when underlying structures have an inherent orientation. In such systems, a tracer may display a different scaling behavior along each spatial dimension and it is necessary to consider the interdependence among the different components. For example, anisotropic scalings emerge in the dynamics of proteins on the cell surface due to the presence of stress fibers [49] and actin filaments [46], ions in the brain [56, 12], macromolecules in the nucleocytoplasm during cell division [41], and species in rocks [55]. Beyond physical systems, financial markets often exhibit non-uniform behavior, prompting the use of multivariate models to capture the joint dynamics of multiple assets or economic indicators [4]. These models help in risk management, option pricing, and forecasting [58, 61]. Multidimensional models that allow dependencies between components have been covered in the literature. Notable examples include multidimensional Brownian motion[4, 45, 54, 3, 22, 19] and multidimensional Ornstein-Uhlenbeck processes[3, 53]. Previous discussions also extended to discrete-time models with dependent components, such as vector autoregressive time series[32, 38, 35, 51], or multidimensional generalized autoregressive conditional heteroskedasticity (GARCH) models [29, 60, 7].
In this article, we propose the construction of a two-dimensional fractional Brownian motion (2D fBm) by introducing a dependence on the underlying noise in the time representation of the process. This construction explicitly considers cross-dependencies and anisotropic scaling by using a matrix-valued Hurst operator that allows direction-dependent scaling properties. We examine the theoretical properties of the proposed construction and derive the auto- and cross-covariance structure of the process and its increments. In addition, we calculate the power spectral density (PSD) of these processes [40], a fundamental measure often used in the characterization of time series across different disciplines [18, 23, 13, 15, 5]. At last, we present numerical simulations that showcase and validate our analytical results. Although multidimensional fBm with dependent components has been investigated in the existing literature[1, 32], our model has the advantage of having a relatively simple construction, based on the correlated Brownian motions used in its integral representation. The new approach uniquely allows for any range of model parameters, a significant advantage over the previous constructions[1]. Consequently, this simplified construction enables a thorough examination of the process characteristics, particularly with respect to its auto- and cross-covariance structures in both the time and frequency domains.
The remainder of the paper is organized as follows. In Section II, we introduce the 2D fBm and discuss its two versions, namely causal and well-balanced. Next, in Section III, we analyze the covariance structure of 2D fBm and its increments, demonstrating the differences for causal and well-balanced cases. In Section IV, we analyze the covariance structure of 2D fBm and the corresponding increments in the frequency domain. In Section V we present numerical simulations, while the last section contains conclusions. In the Appendix, we present additional calculations and plots.
II Model
In this section, we propose two specific constructions of the 2D fBm, which serve as a mathematical tool for modeling anomalous diffusion processes in multidimensional spaces. Unlike the classical one-dimensional case, which relies on scalar parameters to characterize diffusion behavior, we employ a matrix-valued $H$ to capture directional dependencies and cross-dimensional correlations. While we only consider the two-dimensional case, the extension to higher dimensions is straightforward. First, following Stoev and Taqqu[52], let us introduce the following function
$f_{\pm}(x;t,\beta)=(t-x)_{\pm}^{\beta}-(-x)_{\pm}^{\beta},$ (1)
where the notation $(x)_{+}$ denotes the positive part of $x$, i.e.,
$(x)_{+}\equiv\begin{cases}x,&\textrm{if }x>0,\\ 0,&\textrm{if }x\leq 0,\end{cases}$ (2)
and $(x)_{-}$ denotes the negative part of $x$, i.e., $(x)_{-}=(-x)_{+}$. Let us consider the two-dimensional process $\mathbf{X}(t)=[X_{1}(t),X_{2}(t)]^{\prime},t\geq 0$, defined as follows
$\begin{cases}X_{1}(t)&=\sigma_{1}a_{H_{1}}\displaystyle\int_{-\infty}^{\infty}f_{+}\left(s;t,H_{1}-1/2\right)\mathrm{d}\tilde{W}_{1}(s),\\ X_{2}(t)&=\sigma_{2}a_{H_{2}}\displaystyle\int_{-\infty}^{\infty}f_{+}\left(s;t,H_{2}-1/2\right)\mathrm{d}\tilde{W}_{2}(s).\end{cases}$ (3)
A similar alternative $\mathbf{X}^{*}(t)=[X_{1}^{*}(t),X_{2}^{*}(t)]^{\prime},t\geq 0$, utilizing both $f_{+}$ and $f_{-}$ is given by
$\begin{cases}X_{1}^{*}(t)&=\sigma_{1}a^{*}_{H_{1}}\displaystyle\int_{-\infty}^{\infty}\left(f_{+}\left(s;t,H_{1}-1/2\right)+f_{-}(s;t,H_{1}-1/2)\right)\mathrm{d}\tilde{W}_{1}(s),\\ X_{2}^{*}(t)&=\sigma_{2}a^{*}_{H_{2}}\displaystyle\int_{-\infty}^{\infty}\left(f_{+}\left(s;t,H_{2}-1/2\right)+f_{-}(s;t,H_{2}-1/2)\right)\mathrm{d}\tilde{W}_{2}(s).\end{cases}$ (4)
In both cases, $\mathrm{d}\tilde{W}_{1}(t),\mathrm{d}\tilde{W}_{2}(t)$ are $\rho$-correlated Gaussian noises on the real line, $H_{1},H_{2}\in(0,1)$, and $\sigma_{1},\sigma_{2}>0$. The constants $a_{H_{1}}$, $a_{H_{2}}$, $a^{*}_{H_{1}}$, and $a^{*}_{H_{2}}$ are chosen in such a way that the processes $X_{j}(t)$ and $X^{*}_{j}(t)$ for $j=1,2$ have variances $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$ for $t=1$, respectively, that is, $\langle X_{j}^{2}(1)\rangle=\langle X^{*2}_{j}(1)\rangle=\sigma_{j}^{2},\textrm{ for }j=1,2$. Parameter $\rho$ will be referred to as correlation coefficient of the underlying noise. Let us note that for both processes $\mathbf{X}(t)$ and $\mathbf{X}^{*}(t)$, the marginals $X_{j}(t)$ and $X_{j}^{*}(t)$ ($j=1,2$) are fractional Brownian motions with corresponding Hurst parameters $H_{j}$ (cf. [30] for $\mathbf{X}(t)$ and [52] for $\mathbf{X}^{*}(t)$). The difference between the seemingly similar processes $\mathbf{X}(t)$ and $\mathbf{X}^{*}(t)$ lies in the cross-dependence structure between their components, specifically between $X_{1}(t)$ and $X_{2}(t)$, and between $X_{1}^{*}(t)$ and $X_{2}^{*}(t)$, respectively, which we discuss later in this section.
To establish the connection between the processes $\mathbf{X}(t)$ and $\mathbf{X}^{*}(t)$ and the vector fractional Brownian motion discussed in the literature [1, 10, 42], let us consider two independent Brownian motions $W_{1}(t),W_{2}(t)$ on the real line and define $\tilde{W}_{1}(t)$ and $\tilde{W}_{2}(t)$ in the following way
$\begin{cases}\tilde{W}_{1}(t)=W_{1}(t)\\ \tilde{W}_{2}(t)=\rho W_{1}(t)+\sqrt{1-\rho^{2}}W_{2}(t).\end{cases}$ (5)
One can easily show that such a construction leads to $\langle\mathrm{d}\tilde{W}_{1}(t)\mathrm{d}\tilde{W}_{2}(t)\rangle=\rho\mathrm{d}t$. Using the construction from Eq. (5), the processes defined in Eqs. (3) and (4) can be represented more concisely, as shown in the following definitions. Note that the adjectives “causal” and “well-balanced” in the processes’ names are already established in the literature. We follow this convention throughout the article.
Defintition 1 (Causal 2D fBm).
Let $H_{1},H_{2}\in(0,1)$ and $|\rho|\leq 1$. Causal two-dimensional fractional Brownian motion $\mathbf{X}(t),t\geq 0,$ is defined via the following integral representation
$\mathbf{X}(t)=\int_{-\infty}^{\infty}\begin{bmatrix}\sigma_{1}a_{H_{1}}f_{+}(t;s,H_{1}-1/2)&0\\ 0&\sigma_{2}a_{H_{2}}f_{+}(t;s,H_{2}-1/2)\end{bmatrix}\begin{bmatrix}1&0\\ \rho&\sqrt{1-\rho^{2}}\end{bmatrix}\begin{bmatrix}\mathrm{d}W_{1}(s)\\ \mathrm{d}W_{2}(s)\end{bmatrix}$, (6)
where $\sigma_{1},\sigma_{2}>0$. The constants $a_{H_{j}},j=1,2$ are non-negative, and are chosen in such a way that the variances of the marginals $X_{j}(t)$ at time $t=1$ are equal to $\sigma_{j}^{2}$, i.e.,
$a^{2}_{H_{j}}=\frac{\Gamma(2H_{j}+1)\sin(H_{j}\pi)}{\Gamma^{2}\left(H_{j}+\frac{1}{2}\right)}.$ (7)
The process $\mathbf{X}(t)$ is a zero-mean Gaussian process and has a strong connection to the so-called operator fractional Brownian motion[42]. Such relation allows for analysing the covariance structure of the process $\mathbf{X}(t)$ (see Theorem 1 in Section III).
Remark 1.
Let us notice that the marginals $X_{j}(t)$, $j=1,2$ of the causal 2D fBm follow the well-known Mandelbrot and van Ness definition[30] of fBm, i.e.,
$X_{j}(t)=\sigma_{j}a_{H_{j}}\int_{-\infty}^{t}\left((t-s)_{+}^{H_{j}-\frac{1}{2}}-(-s)_{+}^{H_{j}-\frac{1}{2}}\right)\mathrm{d}\tilde{W}_{j}(s),\quad\mathrm{for}\ j=1,2.$ (8)
Contrary to the causal construction presented in Definition 1, the well-balanced case adds to the previous case an anti-causal filtering of a Brownian motion. We now introduce such a process, together with the calculation of the normalization constant (Theorem 6 in Appendix B). The covariance structure of well-balanced 2D fBm is provided in Section III.
Defintition 2 (Well-balanced 2D fBm).
Let $H_{1},H_{2}\in(0,1)$ and $|\rho|\leq 1$. Well-balanced two-dimensional fractional Brownian motion $\mathbf{X}^{*}(t)$ is given by the following representation
$\mathbf{X^{*}}(t)=\int_{-\infty}^{\infty}\begin{bmatrix}g_{1}(s;t)&0\\ 0&g_{2}(s;t)\end{bmatrix}\begin{bmatrix}1&0\\ \rho&\sqrt{1-\rho^{2}}\end{bmatrix}\begin{bmatrix}\mathrm{d}W_{1}(s)\\ \mathrm{d}W_{2}(s)\end{bmatrix}$, (9)
where
$g_{j}(s;t)=\sigma_{j}a^{*}_{H_{j}}\left(f_{+}(s;t,H_{j}-1/2)+f_{-}(s;t,H_{j}-1/2)\right)$ (10)
for $s\in\mathbb{R},t\geq 0,j=1,2$, $\sigma_{1},\sigma_{2}>0$, and the constants $a^{*}_{H_{j}}$’s are given by
$a^{*2}_{H}$ $=\frac{2H(1-2H)\pi}{8\Gamma(2-2H)\cos(H\pi)\Gamma^{2}\left(H+1/2\right)\cos^{2}\left(\frac{\pi(H-1/2)}{2}\right)},$ (11)
to ensure the variances of the marginals $X_{j}^{*}(t)$ at time $t=1$ are equal to $\sigma_{j}^{2}$.
Similarly to the causal 2D fBm, the well-balanced 2D fBm is a zero-mean Gaussian process.
III Covariance structure of 2D fBm
In this section, we discuss the covariance structure of both causal and well-balanced 2D fBms, along with that of their increments.
III.1 Covariance structure of the process
As mentioned, both causal and well-balanced 2D fBm are Gaussian processes; therefore, their covariance structure provides a complete characterization. In this section, we discuss this characterization in time domain.
Theorem 1.
Let $H_{1},H_{2}\in(0,1)$ and $|\rho|\leq 1$. The covariance structure of 2D fBm $\mathbf{Z}(t)=[Z_{1}(t),Z_{2}(t)]^{\prime}$, $t\geq 0$ is as follows
$\gamma_{jk}(t,s)\equiv\langle Z_{j}(t)Z_{k}(s)\rangle=\frac{\sigma_{j}\sigma_{k}}{2}\left(w_{jk}(t)|t|^{H_{j}+H_{k}}+w_{jk}(-s)|s|^{H_{j}+H_{k}}-w_{jk}(t-s)|t-s|^{H_{j}+H_{k}}\right)$ (12)
for $t,s\geq 0$, where $\sigma_{j}^{2}=\langle Z_{j}^{2}(1)\rangle$ and
$w_{jk}(u)=\begin{cases}\rho_{jk}-\eta_{jk}\mathrm{sign}(u),&\quad H_{j}+H_{k}\neq 1,\\ \rho_{jk}-\eta_{jk}\mathrm{sign}(u)\log|u|,&\quad H_{j}+H_{k}=1.\end{cases}$ (13)
The, so called, cross-correlation parameters $\rho_{12}$ and $\rho_{21}$ are given by
$\rho_{12}$ $=\rho_{21}=\rho\frac{\sqrt{\Gamma(2H_{1}+1)\Gamma(2H_{2}+1)\sin(H_{1}\pi)\sin(H_{2}\pi)}}{\Gamma(H_{1}+H_{2}+1)\sin\left(\frac{H_{1}+H_{2}}{2}\pi\right)}\cos\left[\frac{(H_{2}-H_{1})\pi}{2}\right],$ (14)
and $\rho_{11}=\rho_{22}=1$ while the asymmetry parameters $\eta_{jk},j,k=1,2$, depend on the choice of the model, and in the causal 2D fBm (case $\mathbf{Z}(t)=\mathbf{X}(t)$) they are equal to
$\eta_{12}$ $=-\eta_{21}=\rho\frac{\sqrt{\Gamma(2H_{1}+1)\Gamma(2H_{2}+1)\sin(H_{1}\pi)\sin(H_{2}\pi)}}{\Gamma(H_{1}+H_{2}+1)\cos\left(\frac{H_{1}+H_{2}}{2}\pi\right)}\sin\left[\frac{(H_{2}-H_{1})\pi}{2}\right],$ (15)
and $\eta_{11}=\eta_{22}=0$, while for well-balanced 2D fBm (case $\mathbf{Z}(t)=\mathbf{X}(t)^{*}$) we have
$\eta_{11}=\eta_{22}=\eta_{12}=\eta_{21}=0.$ (16)
Proof.
The proof of this theorem is presented in Appendix A and Appendix C. ∎
Remark 2.
Since for well-balanced 2D fBm $\eta_{jk}=0$ for $j,k=1,2$, the covariance function simplifies remarkably. Similar to one-dimensional fBm, it is given by
$\gamma_{jk}(t,s)=\frac{\sigma_{j}\sigma_{k}\rho_{jk}}{2}\left(|t|^{H_{j}+H_{k}}+|s|^{H_{j}+H_{k}}-|t-s|^{H_{j}+H_{k}}\right).$ (17)
Remark 3.
Let us note that for $H_{1}=H_{2}$ both causal and well-balanced 2D fBms define the same process. For $H_{1}=H_{2}=0.5$, both constructions lead to 2D Brownian motion with $\rho$ correlated coordinates[4].
Remark 4.
Let us observe that $\rho_{12}$ given in Eq. (14) can be expressed as
$\rho_{12}=\langle Z_{1}(1)Z_{2}(1)\rangle/\sqrt{\langle Z_{1}^{2}(1)\rangle\langle Z_{2}^{2}(1)\rangle},$ (18)
that is, it plays the role of the cross-correlation coefficient of $\mathbf{Z}(t)$ at time $t=1$.
Theorem 1 highlights the primary distinction between the two considered models. For the causal 2D fBm, we observe that when $\eta_{12}\neq 0$, the cross-covariance function $\gamma_{12}(s,t)$ is asymmetric, meaning $\gamma_{12}(s,t)\neq\gamma_{12}(t,s)$. This is in contrast to the marginal processes, for which $\gamma_{jj}(s,t)=\gamma_{jj}(t,s)$ for any suitable choice of distinct times $t$ and $s$. Conversely, for well-balanced 2D fBm, the cross-covariance function is symmetric (because $\eta_{12}=\eta_{21}=0$).
The correspondence between the correlation coefficient of the underlying noise $\rho$ and the cross-correlation of the process, $\rho_{12}$, presented in Eq. (14), is shown in Figure 1(a) for different Hurst parameter pairs $(H_{1},H_{2})$ of the coordinates $X_{1}(t)$ and $X_{2}(t)$. Let us highlight the fact that, when the difference between $H_{1}$ and $H_{2}$ is large (e.g., $H_{1}=0.2,H_{2}=0.7$, yellow line in the figure) the resulting cross-correlation range is much smaller than the correlation of the underlying noise. In the extreme case of $|\rho|=1$, the parameter $|\rho_{12}|\approx 0.5$. When the difference between $H_{1}$ and $H_{2}$ is small (e.g., $|H_{1}-H_{2}|=0.25\pm 0.05$, orange and violet lines in the figure), the resulting $\rho_{12}$ is similar to the correlation $\rho$.
The correspondence between $\rho$ and the asymmetry parameter $\eta_{12}$, as given in Eq. (15), is presented in Figure 1(b). We observe that $\eta_{12}$ increases with the difference between $H_{1}$ and $H_{2}$, while $\rho_{12}$ decreases.
[…]
Figure 1: The dependence between correlation $\rho$ of the underlying noise and (a) cross-correlation coefficient $\rho_{12}$ and (b) asymmetry parameter $\eta_{12}$ of the process $\mathbf{X}(t)$ and $\mathbf{X}^{*}(t)$. Different solid lines correspond to different Hurst exponents of the coordinates, while the black dashed line corresponds to the identity cases $\rho_{12}=\rho$ or $\eta_{12}=\rho$, shown as a guide to the eye.
Remark 5.
Under the presented construction, the introduced processes $\mathbf{X}(t)$ and $\mathbf{X^{*}}(t)$ are well-defined for all sets of parameters $H_{1},H_{2}\in(0,1)$ and $|\rho|\leq 1$. Such a natural construction ensures that the condition given in Remark 8 in [1] is true. Moreover, we see that $\rho_{12}^{2}\leq\rho^{2}$.
III.2 Covariance structure of increments
For any process $\mathbf{Z}(t)$, the increment process over a step $\delta>0$ is defined as
$\Delta^{\delta}\mathbf{Z}(t)\equiv\mathbf{Z}(t+\delta)-\mathbf{Z}(t).$ (19)
In the following theorem, we present the covariance structure for the increment process of the causal and well-balanced 2D fBm.
Theorem 2.
Let $H_{1},H_{2}\in(0,1)$ and $|\rho|\leq 1$. The covariance structure of the increment process $\Delta^{\delta}\mathbf{Z}(t)$ of causal or well-balanced 2D fBm is as follows
$\gamma_{jk}^{\Delta}(t,s)$ $=\langle\Delta^{\delta}Z_{j}(t)\Delta^{\delta}Z_{k}(s)\rangle=\frac{\sigma_{j}\sigma_{k}\rho_{jk}}{2}\left[w_{jk}(t-s+\delta)|t-s+\delta|^{H_{j}+H_{k}}\right$.
$\qquad+\left.w_{jk}(t-s-\delta)|t-s-\delta|^{H_{j}+H_{k}}-2w_{jk}(t-s)|t-s|^{H_{j}+H_{k}}\right],$ (20)
where the function $w_{jk}$ is given in Eq.(13) and depends on whether we consider causal or well-balanced 2D fBm. The coefficients $\rho_{jk}$ and $\eta_{jk}$ are defined in Theorem 1.
Proof.
The proof of this theorem is a natural consequence of the covariance structure of the introduced processes $\mathbf{X}(t)$ and $\mathbf{X}^{*}(t)$. For the increment process of the causal 2D fBm $\mathbf{X}(t)$ we have
$\langle\Delta^{\delta}X_{j}(t)\Delta^{\delta}X_{k}(s)\rangle$ $=\langle(X_{j}(t+\delta)-X_{j}(t))(X_{k}(s+\delta)-X_{k}(s))\rangle$
$=\gamma_{jk}(t+\delta,s+\delta)-\gamma_{jk}(t,s+\delta)-\gamma_{jk}(t+\delta,s)+\gamma_{jk}(t,s),\quad j,k=1,2.$ (21)
Substituting the covariance function $\gamma_{jk}$ using the results of Theorem 1 yields the desired covariance structure of the increment process. The proof for well-balanced 2D fBm is analogous. ∎
Let us note that for both considered processes, the covariance structure depends only on h=tsh=t-s and the length of the interval $\delta$ on which the increments are taken. For $j=k$, we retain the autocovariance function of fractional Brownian motion corresponding to that coordinate[6], i.e.,
$\langle\Delta^{\delta}X_{j}(t)\Delta^{\delta}X_{j}(s)\rangle=\langle\Delta^{\delta}X^{*}_{j}(t)\Delta^{\delta}X^{*}_{j}(s)\rangle$ $=\frac{\sigma_{j}^{2}}{2}\left(|t-s+\delta|^{2H_{j}}+|t-s-\delta|^{2H_{j}}-2|t-s|^{2H_{j}}\right).$ (22)
IV Spectral Content
The PSD matrix plays a central role in the analysis of multidimensional stochastic processes. This quantity generalizes the concept of a PSD to vector-valued processes. It is particularly valuable in multiple fields, including econometrics [17], structural engineering [11, 21], neuroscience [57], and signal processing [2], where complex systems exhibit interactions across multiple channels or variables. Since PSD is more frequently employed in the context of stationary processes, we begin this section by analyzing the increments of 2D fBm.
IV.1 Spectral content of the increments
In the multivariate setting, for any second-order stationary process $\Delta\mathbf{Z}(t)$, $t\geq 0,$ we can consider the power spectral density matrix $S_{\Delta\mathbf{Z}}(f)$ of a single realization
$S_{\Delta\mathbf{Z}}(f)=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}\int_{0}^{T}e^{i(t-s)f}\Delta\mathbf{Z}(t)\Delta\mathbf{Z}(s)^{\prime}\,\mathrm{d}t\mathrm{d}s,\quad f\in\mathbb{R},$ (23)
with ensemble average
$\langle S_{\Delta\mathbf{Z}}(f)\rangle=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}\int_{0}^{T}e^{i(t-s)f}\gamma_{\Delta\mathbf{Z}}(t-s)\,\mathrm{d}t\mathrm{d}s,\quad f\in\mathbb{R},$ (24)
where $\gamma_{\Delta\mathbf{Z}}(n)=\langle\Delta\mathbf{Z}(n)\Delta\mathbf{Z}(0)^{\prime}\rangle$ for the zero-mean process $\Delta\mathbf{Z}(t)$. Eq. (24) simplifies to
$\langle S_{\Delta\mathbf{Z}}(f)\rangle=\int_{-\infty}^{\infty}e^{itf}\gamma_{\Delta\mathbf{Z}}(t)\,\mathrm{d}t,\quad f\in\mathbb{R},$ (25)
which is known as Wiener-Khinchin theorem[27]. Let us note, that for a $d$-dimensional process, the function $S_{\Delta\mathbf{Z}}$ is a $d\times d$ matrix with the diagonal elements corresponding to marginals and off-diagonal elements describing the, so called, cross power spectral density . To streamline the notation, we will use now $S^{\Delta}$ whenever PSD relates to the increments process, instead of using $S_{\Delta\mathbf{Z}}$.
One can also consider the components of a power spectral density matrix
$\langle S^{\Delta}_{jk}(f)\rangle=\int_{-\infty}^{\infty}e^{itf}\gamma^{\Delta}_{jk}(t)\,\mathrm{d}t,\quad f\in\mathbb{R},~~j,k=1,2.$ (26)
It is worth noting that unlike the one-dimensional version of spectral density, the cross components $S^{\Delta}_{jk}$ for $j\neq k$ might not be real-valued. In general, it is true that $S^{\Delta}_{jk}(f)=\overline{S^{\Delta}_{kj}(f)}$ for all $f$’s, where $\overline{z}$ denotes the complex conjugate of $z$.
Theorem 3.
Let $H_{1},H_{2}\in(0,1)$, $H=\mathrm{diag}(H_{1},H_{2})$ be a diagonal matrix with elements $H_{1},H_{2}$, and $|\rho|\leq 1$. Additionally, let matrices $C_{\textrm{c}}=[c^{\textrm{c}}_{jk}]_{j,k=1,2}$ and $C_{\textrm{wb}}=[c^{\textrm{wb}}_{jk}]_{j,k=1,2}$ that correspond to causal and well-balanced 2D fBm, respectively, have the following elements
$c^{\textrm{c}}_{jj}$ $=\frac{1}{2\pi}\sigma^{2}_{j}\Gamma^{2}(H_{j}+0.5)a^{2}_{H_{j}},$ (27)
$c^{\textrm{c}}_{jk}$ $=\frac{1}{2\pi}\rho\sigma_{j}\sigma_{k}\Gamma(H_{j}+0.5)\Gamma(H_{k}+0.5)a_{H_{j}}a_{H_{k}}e^{-i\frac{\pi}{2}(H_{j}-H_{k})},\quad j\neq k,$ (28)
$c^{\textrm{wb}}_{jj}$ $=\frac{2}{\pi}\sigma^{2}_{j}\cos^{2}\left(\frac{(H_{j}-0.5)\pi}{2}\right)\Gamma^{2}(H_{j}+0.5)a^{*2}_{H_{j}},$ (29)
$c^{\textrm{wb}}_{jk}$ $=\frac{1}{2\pi}\rho\sigma_{j}\sigma_{k}\cos\left(\frac{(H_{j}-0.5)\pi}{2}\right)\cos\left(\frac{(H_{k}-0.5)\pi}{2}\right)\Gamma(H_{j}+0.5)\Gamma(H_{k}+0.5)a^{*}_{H_{j}}a^{*}_{H_{k}},\quad j\neq k,$ (30)
for $j,k=1,2$. Constants $a_{H_{j}}$ and $a^{*}_{H_{j}}$ are given in Eqs. (7) and (11), respectively. Then, for the increments of 2D fBm the power spectral density matrix $S^{\Delta}(f)$ is given by
$\langle S^{\Delta}(f)\rangle$ $=\left|1-e^{-if}\right|^{2}\sum_{n=-\infty}^{\infty}\left[(f+2\pi n)_{+}^{-D}\tilde{C}(f+2\pi n)_{+}^{-D}+(f+2\pi n)_{-}^{-D}\overline{\tilde{C}}(f+2\pi n)_{-}^{-D}\right]/(f+2\pi n)^{2},$ (31)
where $(x)_{+}\equiv\max\{x,0\},(x)_{-}\equiv\max\{-x,0\},D=H-0.5I_{2}$, and $\overline{C}$ denotes the element-wise complex conjugate of matrix $C$. Matrix $\tilde{C}$ is equal to $C_{\textrm{c}}$ or $C_{\textrm{wb}}$ depending if we consider causal or well-balanced 2D fBm. The components $\langle S^{\Delta}_{jk}\rangle$ can be thus expressed as
$\langle S^{\Delta}_{jk}(f)\rangle=|1-e^{-if}|^{2}\sum_{n=-\infty}^{\infty}\left[(f+2\pi n)_{+}^{1-H_{j}-H_{k}}\tilde{c}_{jk}+(f+2\pi n)_{-}^{1-H_{j}-H_{k}}\overline{\tilde{c}}_{jk}\right]/(f+2\pi n)^{2},$ (32)
where $\tilde{c}_{jk}$ are elements of the appropriate matrix, $C_{\textrm{c}}$ for the causal case or $C_{\textrm{wb}}$ for the well-balanced one. Moreover, for $f\to 0$ we have
$\langle S^{\Delta}(f)\rangle\sim f^{-D}\tilde{C}f^{-D}.$ (33)
The element-wise asymptotic is as follows
$\langle S^{\Delta}_{jk}(f)\rangle\sim\tilde{c}_{jk}f^{-(d_{j}+d_{k})}=\tilde{c}_{jk}f^{-(H_{j}+H_{k}-1)}.$ (34)
Proof.
The proof of this theorem is presented in Appendix D. ∎
Remark 6.
Matrices $C_{\textrm{c}}$ and $C_{\textrm{wb}}$ given in Theorem 3 have real elements on the diagonal, thus, $\langle S^{\Delta}_{jj}\rangle,j=1,2$, is also a real (and non-negative) function.
Remark 7.
The direct derivation of matrix $C$ given in Theorem 3 be found in the proof of Theorem 1 given in Appendix A and Appendix C.
IV.2 Spectral content of the process
An extension of PSD calculated on the stationary process (here, for the increments process of 2D fBm), one considers PSD calculated based on the information from the trajectory. In particular, given a nonstationary 1D process $Z(t)$, $0\leq t\leq T$, measured over a time $T$, the PSD is typically defined as [40, 39, 26, 25]
$S_{Z}(f,T)\equiv\frac{1}{T}\left|\int_{0}^{T}e^{ift}Z(t)\,\mathrm{d}t\right|^{2}.$ (35)
Note that in contrast to Eq. (23), the PSD depends on measurement time $T$. In $d$ dimensions, the natural extension of the PSD defined in Eq. (35) for a real process $\mathbf{Z}(t)$, $0\leq t\leq T$, can be written in an equivalent form,
$S_{\mathbf{Z}}(f,T)=\frac{1}{T}\int_{0}^{T}e^{ift}\mathbf{Z}(t)\,\mathrm{d}t\left(\int_{0}^{T}e^{ifs}\mathbf{Z}(s)\,\mathrm{d}s\right)^{*},$ (36)
where $(\mathbf{x})^{*}$ is the Hermitian transpose (also known as the conjugate transpose) of $\mathbf{x}$, i.e., $(\mathbf{x})^{*}=\overline{\mathbf{x}}^{\prime}$. Again, the function $S_{\mathbf{Z}}$ is a $d\times d$ matrix with the diagonal elements corresponding to marginals, and off-diagonal elements corresponding to the cross power spectral density. Eq. (36) can be written using a double integral
$S_{\mathbf{Z}}(f,T)=\frac{1}{T}\int_{0}^{T}\int_{0}^{T}e^{if(t-s)}\mathbf{Z}(t)\mathbf{Z}^{\prime}(s)\mathrm{d}s\,\mathrm{d}t.$ (37)
Taking the expected value, we can also calculate the ensemble-averaged PSD
$\langle S_{\mathbf{Z}}(f,T)\rangle\equiv\frac{1}{T}\int_{0}^{T}\int_{0}^{T}e^{if(t-s)}\langle\mathbf{Z}(t)\mathbf{Z}^{\prime}(s)\rangle\mathrm{d}s\,\mathrm{d}t,$ (38)
alternatively, written component-wise it is
$\langle S_{\mathbf{Z},jk}(f,T)\rangle\equiv\frac{1}{T}\int_{0}^{T}\int_{0}^{T}e^{if(t-s)}\langle Z_{j}(t)Z_{k}(s)\rangle\mathrm{d}s\,\mathrm{d}t,~~j,k=1,2.$ (39)
Let us note that for $j=k$ we obtain the classical PSD given in Eq. (35) of the 1D process corresponding to the marginals. Here, we further use the simplified notation $S_{jk}$ instead of $S_{\mathbf{Z},jk}$ to streamline the notation whenever it is clear which process is referred to. In both cases (causal and well-balanced), ensemble-averaged PSD for 2D fBm can be expressed as
$\langle S_{jk}(\tilde{\omega},T)\rangle=T^{H_{j}+H_{k}+1}\frac{\sigma_{j}\sigma_{k}}{2}\int_{0}^{1}\int_{0}^{1}e^{i\tilde{\omega}(t-s)}\left(w_{jk}(x)x^{H_{j}+H_{k}}+w_{jk}(-y)y^{H_{j}+H_{k}}-w_{jk}(x-y)|x-y|^{H_{j}+H_{k}}\right)\mathrm{d}x\,\mathrm{d}y,$ (40)
where $\tilde{\omega}=fT$. The function $w_{jk}$ is defined in Eq. (13) with parameters $\rho_{jk}$ and asymmetry parameters $\eta_{jk}$ defined in Theorem 1. In the following theorem, we present the behavior of such a function.
Theorem 4.
For the 2D fBm with $H_{1}+H_{2}\neq 1$, the ensemble-averaged PSD for coordinates $j,k=1,2$ has the form
$\langle S_{jk}(\tilde{\omega},T)\rangle=T^{H_{j}+H_{k}+1}\sigma_{j}\sigma_{k}$ $\left\{\rho_{jk}\left[\frac{1-\cos\tilde{\omega}}{\tilde{\omega}}\mathcal{S}_{jk}-\left(1-\frac{\sin\tilde{\omega}}{\tilde{\omega}}\right)\mathcal{C}_{jk}+\frac{\mathrm{d}}{\mathrm{d}\tilde{\omega}}\mathcal{S}_{jk}\right]\right$.
$+$ $\eta_{jk}\left.i\left[\frac{1-\cos\tilde{\omega}}{\tilde{\omega}}\mathcal{C}_{jk}-\left(1+\frac{\sin\tilde{\omega}}{\tilde{\omega}}\right)\mathcal{S}_{jk}-\frac{\mathrm{d}}{\mathrm{d}\tilde{\omega}}\mathcal{C}_{jk}\right]\right\},$ (41)
where
[…]
Parameters $\rho_{jk}$ and $\eta_{jk}$ are given in Theorem 1.
Proof.
The proof of this theorem is included in Appendix E. ∎
Remark 8.
We can also use alternative forms for functions $\mathcal{C}$ and $\mathcal{S}$ defined in Eqs. (42) and (43), respectively, to obtain the expressions without derivatives. Alternatively, using integration by parts and the identities
$\frac{\mathrm{d}}{\mathrm{d}\tilde{\omega}}\mathcal{S}_{jk}$ $=\frac{\sin\tilde{\omega}}{\tilde{\omega}}-\frac{H_{j}+H_{k}+1}{\tilde{\omega}}\mathcal{S}_{jk},$ (44)
$\frac{\mathrm{d}}{\mathrm{d}\tilde{\omega}}\mathcal{C}_{jk}$ $=\frac{\cos\tilde{\omega}}{\tilde{\omega}}-\frac{H_{j}+H_{k}+1}{\tilde{\omega}}\mathcal{C}_{jk},$ (45)
we obtain the equivalent expression
[…]
For marginals (i.e., when $j=k$), the expression reduces to the known formula for PSD of 1-dimensional fBm with Hurst parameter $H=H_{j}=H_{k}$ [25], since $\eta_{jj}=0$. In contrast, for $j\neq k$, the cross power spectral density acquires an imaginary part in the causal case, reflecting time-asymmetric dependence between components. This imaginary component vanishes in the well-balanced case, where $\eta_{jk}\equiv 0$ (cf. Eq. (16)).
Theorem 5 (Asymptotic behavior of the ensemble-averaged PSD).
The ensemble-averaged PSD $\langle S_{jk}(\tilde{\omega},T)\rangle$ of the 2D fBm admits the following asymptotic regimes as $\tilde{\omega}\to\infty$ if $H_{j}+H_{k}\neq 1$ or for well-balanced 2D fBm
$\Re\langle S_{jk}(\tilde{\omega},T)$ $\rangle\sim\quad T^{H_{j}+H_{k}+1}\sigma_{j}\sigma_{k}\rho_{jk}\left\{\left[\frac{1}{\tilde{\omega}^{2}}-\frac{(H_{j}+H_{k})\sin\tilde{\omega}}{\tilde{\omega}^{3}}+O\left(\frac{1}{\tilde{\omega}^{4}}\right)\right]\right$.
$\left.+a_{jk}\left[\frac{1}{\tilde{\omega}^{H_{j}+H_{k}+1}}-\frac{(H_{j}+H_{k}+\cos\tilde{\omega})\cot\left(\frac{\pi}{2}(H_{j}+H_{k})\right)+\sin\tilde{\omega}}{\tilde{\omega}^{H_{j}+H_{k}+2}}\right]\right\},$ (47)
$\Im\langle S_{jk}(\tilde{\omega},T)\rangle$ $\sim\quad T^{H_{j}+H_{k}+1}\sigma_{j}\sigma_{k}\eta_{jk}\left\{\left[\frac{2}{\tilde{\omega}^{2}}-\frac{3-\cos\tilde{\omega}}{\tilde{\omega}^{3}}+O\left(\frac{1}{\tilde{\omega}^{4}}\right)\right]\right$.
$\left.-a_{jk}\left[\frac{\cot\left(\frac{\pi}{2}(H_{j}+H_{k})\right)}{\tilde{\omega}^{H_{j}+H_{k}+1}}+\frac{2+H_{j}+H_{k}-\cos\tilde{\omega}\sin\tilde{\omega}\cot\left(\frac{\pi}{2}(H_{j}+H_{k})\right)}{\tilde{\omega}^{H_{j}+H_{k}+2}}\right]\right\},$ (48)
where $\Re$ and $\Im$ denote the real and imaginary parts, respectively. Parameters $\rho_{jk}$ and $\eta_{jk}$ are given in Theorem 1. In the formulas above, the parameters $a_{jk}=\Gamma(H_{j}+H_{k})\sin\left(\frac{\pi}{2}(H_{j}+H_{k})\right)$.
Proof.
The proof of this theorem is based on the asymptotic expansion of the incomplete gamma function and is presented in details in Appendix F. ∎
The behavior of ensemble-averaged PSD od 2D fBm depends explicitly on the Hurst parameters $H_{j}$, $H_{k}$, as well as the parameters $\rho_{jk}$ and $\eta_{jk}$. In the following remark we present the details.
Remark 9.
The asymptotic behavior of the ensemble-averaged PSD $\langle S_{jk}(\tilde{\omega},T)\rangle$ for 2D fBm depends on whether the sum of Hurst parameters $H_{j}+H_{k}$ is greater than or smaller than 1. In particular, when $H_{j}+H_{k}>1$, we observe dependence on measurement time $T$. More precisely, we have
$\Re\langle S_{jk}(f,T)\rangle\sim\frac{\sigma_{j}\sigma_{k}\rho_{jk}}{f^{2}}T^{H_{j}+H_{k}-1}+\frac{2c_{H}\sigma_{j}\sigma_{k}}{f^{H_{j}+H_{k}+1}}+o(1).$ (49)
Conversely, for $H_{j}+H_{k}<1$, the decay is faster, and the cross-component interactions diminish more rapidly with frequency
$\Re\langle S_{jk}(f,T)\rangle\sim\frac{\sigma_{j}\sigma_{k}\rho_{jk}}{f^{H_{j}+H_{k}+1}}+\frac{2\sigma_{j}\sigma_{k}}{T^{1-H_{j}-H_{k}+1}}+o(T^{H_{j}+H_{k}-1}).$ (50)
This dichotomy highlights the crucial role of $H_{j}+H_{k}$ on the spectral structure of the process.
V Numerical simulations
In this section, we present the comparison of the analytical results obtained in the previous parts of this article with numerical simulations. For the numerical simulations of 2D fBm we apply the Wood and Chan’s algorithm[9], which is based on embedding the covariance matrix of 2D fBm increments into a circulant matrix. All of the presented numerical results are based on the ensemble of $N=5,000$ trajectories of length $T=2^{16}=65,536$ with a time step $1$.
V.1 Trajectories
To visualize 2D fBm and evaluate our analytical results, we performed extensive numerical simulations. Figures 2 and 3 show sample trajectories of the 2D fBm. In both figures, upper row corresponds to the case of independent coordinates, i.e., $\rho=\rho_{12}=0$, while the bottom row corresponds to the case $\rho_{12}=0.5$. In Figure 2 we present the case with $H_{1}=H_{2}$, that is, a case when causal and well-balanced processes coincide. In Figure 3 we focus on the case with $H_{1}\neq H_{2}$ and the trajectories are realizations of the causal 2D fBm. We present more trajectories to highlight the difference in the behavior of the process in the cardinal directions. For example, on panels (b) and (d), the process is characterized by subdiffusion on the horizontal axis (1st coordinate) while on the vertical it is a superdiffusion (2nd coordinate).
[…]
Figure 2: Sample trajectories of causal 2D fBm with $H_{1}=H_{2}=H$. Each panel corresponds to a given set of parameters $(H,\rho_{12})$, where the top row (a-c) corresponds to uncorrelated components and the bottom row (d-f) to cross-correlation $\rho_{12}=0.5$. Pairs of panels in each column have the same $H$. The trajectories are shifted on the horizontal axis for clarity.
[…]
Figure 3: Sample trajectories of causal 2D fBm with different Hurst exponents, $H_{1}\neq H_{2}$. Each panel corresponds to a given set of parameters $(H_{1},H_{2},\rho_{12})$, where the top row (a-c) corresponds to uncorrelated components and the bottom row (d-f) to cross-correlation $\rho_{12}=0.5$. The trajectories are shifted on the horizontal axis for clarity.
V.2 Autocovariance
The different cases of cross-covariance functions $\gamma_{12}$ given in Theorem 1 are presented in Figures 4 and 5. In both of these figures, the cross-covariance function for the causal version of the process is plotted using a solid line, while for well-balanced case, a dashed line is used. Figure 4 showcases the situation when $H_{1}=H_{2}$. However, as discussed in Remark 3, in such a case both causal and well-balanced cases are the same – that is why the continuous and dashed lines overlap each other. We see that the cross-covariance function behaves similarly to that we expect for the fractional Brownian motion with the corresponding $H$. The only difference between cross-covariances of the causal and well-balanced processes is the rescaling of such function by a factor of $\rho_{12}$. Thus, each case (subdiffusive $H_{1}=H_{2}=0.2$ – blue line, left panel; diffusive $H_{1}=H_{2}=0.5$ – orange line, middle panel; superdiffusive $H_{1}=H_{2}=0.7$ – yellow line, right panel) retains the behavior.
A more interesting situation is presented in Figure 5, where the corresponding Hurst parameters $H_{1}$ and $H_{2}$ are not equal. Here, the dashed lines corresponding to the well-balanced case are different from the ones for causal 2D fBm. Particularly, it seems that the ‘strength’ of the cross-covariance is usually smaller (in the absolute sense) in such cases. The most interesting is the case of $H_{1}=0.2,H_{2}=0.7,\rho_{12}=0.5$ (orange line, middle panel) – for which we change the type of memory for lags $h>0$. For the causal case, the cross-covariance function was positive, while the addition of the well-balanced element makes it negative. Another interesting effect can be observed for the causal version for $H_{1}=0.2,H_{2}=0.5$ and $H_{1}=0.5,H_{2}=0.7$ and $\rho_{12}=0.5$ in both cases. The cross-covariance function disappears for $h>0$ and $h<0$, respectively. It is due to the fact that via the choice of parameters we obtained $\rho_{12}-\eta_{12}\mathrm{sign}{h}=0$ thus cancelling every element in the cross-covaraince function for the increments for lags $|h|\geq 1$ (cf. the relation between $\rho_{12}$ and $\eta_{12}$ in Eq. (70) of the Appendix).
[…]
Figure 4: Cross-covariance function depending on $H_{1},H_{2}$ and $\rho_{12}$ for $H_{1}=H_{2}$. (a) $H_{1}=H_{2}=0.2$, (b) $H_{1}=H_{2}=0.5$, and (c) $H_{1}=H_{2}=0.7$. The solid lines represent the causal version of the model and dashed lines the well-balanced case. Markers correspond to the estimated values of the cross-covariance function; circles correspond to the causal case, squares to the well-balanced case. Here, casual and well-balanced cases overlap, as for $H_{1}=H_{2}$ the model is time-reversible regardless of the definition (cf. Eq. (15)).
[…]
Figure 5: Cross-covariance function depending on $H_{1},H_{2}$ and $\rho_{12}$ for $H_{1}\neq H_{2}$. (a) $H_{1}=0.2,H_{2}=0.5$, (b) $H_{1}=0.2,H_{2}=0.7$, and (c) $H_{1}=0.5,H_{2}=0.7$. Solid lines represent the causal version of the model, the dashed lines the well-balanced case. Markers correspond to the estimated values of the cross-covariance function; circles correspond to the causal case, squares to the well-balanced case.
V.3 Spectral content
In Figure 6 we present a comparison between estimated ensemble-averaged (cross-)PSD to their theoretical asymptotics for the 2D fBm. As in previous figures, solid lines represent estimates from simulations, while dashed lines indicate the theoretical asymptotics. Different colored solid lines correspond to different sets of parameters $(H_{1},H_{2})$, whereas $\rho_{12}=0.5$ for all the cases. Panels (a) and (b) present the real part, i.e., $\Re\langle S_{\mathbf{X},12}(f,T)\rangle$, with panel (a) corresponding to the case with $H_{1}=H_{2}$ (i.e., when the causal and well-balanced 2D fBms coincide), while panel (b) represents the causal 2D fBm with $H_{1}\neq H_{2}$. Panel (c) showcases the imaginary part, i.e., $\Im\langle S_{\mathbf{X},12}(f,T)\rangle$ for the sets $H_{1}\neq H_{2}$ corresponding to panel (b) (as for the $H_{1}=H_{2}$ case, the cross-PSD is 0).
In all panels, we see that the asymptotic behavior of the ensemble-averaged cross-PSD depends on the $H_{1}+H_{2}$ value: if it is smaller than 1, the asymptotic decay behaves like $f^{-1-H_{1}-H_{2}}$ (for large $f$), whereas if $H_{1}+H_{2}\geq 1$, then the asymptotic decay is proportional to $f^{-2}$.
[…]
Figure 6: Ensemble-averaged (cross-)PSD. Solid lines correspond to PSD estimated from the simulated trajectories, dashed lines to their asymptotics. (a) $\Re\langle S_{\mathbf{X},12}(f,T)\rangle$, $H_{1}=H_{2}$; (b) $\Re\langle S_{\mathbf{X},12}(f,T)\rangle$, $H_{1}\neq H_{2}$; and (c) $\Im\langle S_{\mathbf{X},12}(f,T)\rangle$, $H_{1}\neq H_{2}$. Since the real part for both causal and well-balanced case is the same, we plot only only one corresponding line to each set of parameteres. Conversely, the imaginary part for the well-balanced case is 0, so we plot only $\Im\langle S_{\mathbf{X},12}\rangle$ for the causal 2D fBm. In all the panels, $T=2^{16}=65,536$ and the lines are based on $N=5,000$ trajectories.
Similarly to Figure 6, in Figure 7 we present a comparison between estimated ensemble-averaged (cross-)PSD to their theoretical asymptotics for the increments of 2D fBm. Again, solid lines represent estimates from simulations, while dashed lines indicate the theoretical asymptotics. Different colored solid lines correspond to different sets of parameters $(H_{1},H_{2})$, whereas $\rho_{12}=0.5$ for all the cases. Panels (a) and (b) present the real part, i.e., $\Re\langle S_{\Delta\mathbf{X},12}(f,T)\rangle$, with panel (a) corresponding to the case with $H_{1}=H_{2}$ (i.e., when the causal and well-balanced 2D fBms coincide), while panel (b) represents the causal 2D fBm increments with $H_{1}\neq H_{2}$. Panel (c) showcases the imaginary part, i.e., $\Im\langle S_{\Delta\mathbf{X},12}(f,T)\rangle$ for the sets $H_{1}\neq H_{2}$ corresponding to panel (b) (as for the $H_{1}=H_{2}$ case, the cross-PSD is 0).
In all panels, we see that the asymptotic behavior of the ensemble-averaged (cross-)PSD for increments behaves in a similar way in all cases, regardless if $H_{1}+H_{2}\gtrless 1$. The asymptotic behavior, as $f\to 0$ is proportional to $f^{1-H_{1}-H_{2}}$. Here, we see a quantitative difference of diverging or converging (cross-)PSD around 0 – if $H_{1}+H_{2}>1$ then the (cross-)PSD diverges (in 1D, it corresponds to one of the definitions of long memory [42]), while for $H_{1}+H_{2}$ the (cross-)PSD converges to 0.
[…]
Figure 7: Ensemble-averaged (cross-)PSD for increments. Continuous lines correspond to PSD estimated from the simulated trajectories’ increments, dashed lines to their asymptotics. (a) $\Re\langle S^{\Delta}_{12}(f,T)\rangle$, $H_{1}=H_{2}$; (b) $\Re\langle S^{\Delta}_{12}(f)\rangle$, $H_{1}\neq H_{2}$; and (c) $\Im\langle S^{\Delta}_{12}(f)\rangle$, $H_{1}\neq H_{2}$. Since the real part for both causal and well-balanced case is the same, we plot only only one corresponding line to each set of parameters. Conversely, the imaginary part for the well-balanced case is 0, so we plot only $\Im\langle S^{\Delta}_{12}(f)\rangle$ for the causal 2D fBm’s increments. In all the panels, $T=2^{16}=65,536$ and the lines are based on $N=5,000$ trajectories.
VI Conclusions
In this work, we introduced a natural construction of 2D fBm that accounts for anisotropic scaling and cross-dimensional dependencies, extending the classical framework of Mandelbrot and van Ness to multivariate settings. By incorporating correlated Gaussian noises and a matrix-valued Hurst operator, we provided two distinct formulations, causal and well-balanced 2D fBm, and derived their respective auto- and cross-covariance across the two components. Our analysis highlights key differences between these two formulations, particularly in the asymmetry of causal fBm and implications for time-reversal. Furthermore, we analyzed the spectral properties of the processes, presenting both full analytical derivations and asymptotic behaviors in the frequency domain. Our results reveal how variations in the Hurst exponents and correlation structure manifest in the spectral content, offering new insights into the interpretation of multivariate time series in diverse systems. An especially interesting case arises when $H_{1}=H_{2}$. In this situation, the process remains a 2D fBm under arbitrary rotations. In contrast, when $H_{1}\neq H_{2}$, a rotation mixes the coordinates in such a way that the resulting marginals are no longer fBm, but instead they exhibit more complex dependencies.
Extensive numerical simulations validated our theoretical findings, demonstrating consistency in both the time and frequency domains. The proposed framework thus provides a robust tool for modeling and analyzing complex anomalous diffusion processes in multidimensional environments, especially in settings where directional dependence and inter-component correlations cannot be ignored. This study opens several avenues for further exploration. A potential immediate extension involves generalizing the proposed constructions to higher dimensions. Our work opens the way to the application of 2D fBm to empirical data in fields such as neuroscience, single-particle tracking, structural dynamics, and financial econometrics, which may yield new insights into underlying spatiotemporal correlations. While all analyses are shown for the two-dimensional case, they can be readily generalized to any dimension.
Acknowledgements
This work was supported by NCN OPUS project No. 2024/53/B/HS4/00433 (to AW) and by the National Science Foundation (NSF) Grant 2102832 (to DK).
Appendix A Proof of Theorem 1, causal case and some additional remarks
Proof.
Comparing the definition of the proposed causal 2D fBm process (cf. Definition 1) to the definition of operator fractional Brownian motion $B_{H}(t),t\geq 0,$ given in formula (9.3.22) from [42],
$B_{H}(t)=\int_{\mathbb{R}}((t-u)_{+}^{D}-(-u)_{+}^{D})M_{+}+((t-u)_{-}^{D}-(-u)_{-}^{D})M_{-})B(\mathrm{d}u),$ (51)
we see that for our causal 2D fBm is equivalent when the matrices $M_{+}$ and $M_{-}$ are as follows
$M_{+}$ $=\begin{bmatrix}\sigma_{1}a_{H_{1}}&0\\ \rho\sigma_{2}a_{H_{2}}&\sqrt{1-\rho^{2}}\sigma_{2}a_{H_{2}}\end{bmatrix}$, (52)
$M_{-}$ $=0\cdot I_{2},$ (53)
for a size-two identity matrix $I_{2}$.
Utilizing (9.3.28-29) from [42], we can calculate $A=A_{1}+iA_{2}$ that later will be used for determining the covariance structure of the proposed process
$A_{1}$ $=\frac{1}{\sqrt{2\pi}}\Gamma(D+I)\cdot\sin\left(\frac{D\pi}{2}\right)M_{+},$ (54)
$A_{2}$ $=\frac{1}{\sqrt{2\pi}}\Gamma(D+I)\cdot\cos\left(\frac{D\pi}{2}\right)M_{+},$ (55)
where $D=H-\frac{1}{2}I$ is a 2x2 matrix, and all of the present functions ($\Gamma,\sin,\cos$) are understood as the primary matrix functions (def. 9.3.4 in [42]). Therefore, we can explicitly write
$A_{1}$ $=\frac{1}{\sqrt{2\pi}}\begin{bmatrix}\Gamma(H_{1}+\frac{1}{2})&0\\ 0&\Gamma(H_{2}+\frac{1}{2})\end{bmatrix}\cdot\begin{bmatrix}\sin\left(\frac{(H_{1}-\frac{1}{2})\pi}{2}\right)&0\\ 0&\sin\left(\frac{(H_{2}-\frac{1}{2})\pi}{2}\right)\end{bmatrix}\cdot\begin{bmatrix}\sigma_{1}a_{H_{1}}&0\\ \rho\sigma_{2}a_{H_{2}}&\sqrt{1-\rho^{2}}\sigma_{2}a_{H_{2}}\end{bmatrix}$
$=\begin{bmatrix}\Gamma(H_{1}+\frac{1}{2})\sin\left(\frac{(H_{1}-\frac{1}{2})\pi}{2}\right)\sigma_{1}a_{H_{1}}&0\\ \rho\Gamma(H_{2}+\frac{1}{2})\sin\left(\frac{(H_{2}-\frac{1}{2})\pi}{2}\right)\sigma_{2}a_{H_{2}}&\sqrt{1-\rho^{2}}\Gamma(H_{2}+\frac{1}{2})\sin\left(\frac{(H_{2}-\frac{1}{2})\pi}{2}\right)\sigma_{2}a_{H_{2}}\end{bmatrix}$, (56)
and similarly for $A_{2}$
$A_{2}$ $=\begin{bmatrix}\Gamma(H_{1}+\frac{1}{2})\cos\left(\frac{(H_{1}-\frac{1}{2})\pi}{2}\right)\sigma_{1}a_{H_{1}}&0\\ \rho\Gamma(H_{2}+\frac{1}{2})\cos\left(\frac{(H_{2}-\frac{1}{2})\pi}{2}\right)\sigma_{2}a_{H_{2}}&\sqrt{1-\rho^{2}}\Gamma(H_{2}+\frac{1}{2})\cos\left(\frac{(H_{2}-\frac{1}{2})\pi}{2}\right)\sigma_{2}a_{H_{2}}\end{bmatrix}$. (57)
Thus, $A=A_{1}+iA_{2}$ can be written as
$A$ $=\begin{bmatrix}\Gamma(H_{1}+\frac{1}{2})\exp\{i\frac{\pi}{2}(\frac{3}{2}-H_{1})\}\sigma_{1}a_{H_{1}}&0\\ \rho\Gamma(H_{2}+\frac{1}{2})\exp\{i\frac{\pi}{2}(\frac{3}{2}-H_{2})\}\sigma_{2}a_{H_{2}}&\sqrt{1-\rho^{2}}\Gamma(H_{2}+\frac{1}{2})\exp\{i\frac{\pi}{2}(\frac{3}{2}-H_{2})\}\sigma_{2}a_{H_{2}}\end{bmatrix}$. (58)
In order to obtain the covariance structure, we calculate $C=[c_{jk}]_{j,k=1,2}=AA^{*}$ (from Proposition 9.3.19 [42]). In our case, its elements are given by
$\begin{cases}c_{jj}&=\frac{1}{2\pi}\sigma^{2}_{j}\Gamma^{2}(H_{j}+0.5)a^{2}_{H_{j}},\\ c_{jk}&=\frac{1}{2\pi}\rho\sigma_{j}\sigma_{k}\Gamma(H_{j}+0.5)\Gamma(H_{k}+0.5)a_{H_{j}}a_{H_{k}}e^{-i\frac{\pi}{2}(H_{j}-H_{k})},\quad j\neq k,\end{cases}$ (59)
for $j,k=1,2$. Note that the matrix $C$ is Hermitian, $C=C^{*}$, and thus its diagonal elements are real. According to Proposition 9.3.19 [42], the covariance structure is
$\gamma_{jk}(t,s)\equiv\langle X_{j}(t)X_{k}(s)\rangle=\frac{\sigma_{j}\sigma_{k}}{2}\left(w_{jk}(t)|t|^{H_{j}+H_{k}}+w_{jk}(-s)|s|^{H_{j}+H_{k}}-w_{jk}(t-s)|t-s|^{H_{j}+H_{k}}\right),$ (60)
where $\sigma_{j}^{2}=\langle X_{j}^{2}(1)\rangle$ and
$w_{jk}(u)=\begin{cases}\rho_{jk}-\eta_{jk}\mathrm{sign}(u),&\quad H_{j}+H_{k}\neq 1,\\ \rho_{jk}-\eta_{jk}\mathrm{sign}(u)\log|u|,&\quad H_{j}+H_{k}=1.\end{cases}$ (61)
The connection between the elements of matrix $C$ given in (59) and the parameters $\rho_{jk}$ and $\eta_{jk}$ are as follows
$\sigma_{j}\sigma_{k}\rho_{jk}=4b_{1}\left(\frac{H_{j}+H_{k}}{2}\right)\Re(c_{jk}),$ (62)
$\sigma_{j}\sigma_{k}\eta_{jk}=4b_{2}\left(\frac{H_{j}+H_{k}}{2}\right)\Im(c_{jk}),$ (63)
where
$b_{1}(H)$ $=\begin{cases}\frac{\Gamma(2-2H)\cos(H\pi)}{2H(1-2H)},&\quad H\neq\frac{1}{2},\\ \frac{\pi}{2},&\quad H=\frac{1}{2},\end{cases}$ (64)
$b_{2}(H)$ $=\frac{\Gamma(2-2H)\sin(H\pi)}{2H(1-2H)}.$ (65)
Let us calculate $\rho_{jk}$ explicitly. First, consider $j=k$, and assume that $H_{j}\neq\frac{1}{2}$. We have
$\sigma_{j}\sigma_{j}\rho_{jj}$ $=4b_{1}\left(\frac{H_{j}+H_{j}}{2}\right)\Re(c_{jj})=4b_{1}(H_{j})\Re(c_{jj})$
$=4\frac{\Gamma(2-2H_{j})\cos(H_{j}\pi)}{2H_{j}(1-2H_{j})}\frac{1}{2\pi}\Gamma^{2}\left(H_{j}+\frac{1}{2}\right)\sigma^{2}_{j}a^{2}_{H_{j}}$
$=4\sigma^{2}_{j}\frac{\overbrace{\Gamma(1-2H_{j})}^{\Gamma(z)\Gamma(1-z)=\frac{\pi}{\sin(\pi z)}}\cos(H_{j}\pi)}{2H_{j}}\frac{1}{2\pi}\Gamma^{2}\left(H_{j}+\frac{1}{2}\right)a^{2}_{H_{j}}$
$=2\sigma_{j}^{2}\frac{\cos(H_{j}\pi)}{\underbrace{2H_{j}\Gamma(2H_{j})}_{\Gamma(2H_{j}+1)}\sin(2H_{j}\pi)}\Gamma^{2}\left(H_{j}+\frac{1}{2}\right)\underbrace{a^{2}_{H_{j}}}_{=\frac{\Gamma(2H_{j}+1)\sin(H_{j}\pi)}{\Gamma^{2}(H_{j}+\frac{1}{2})}}$
$=\sigma_{j}^{2}.$ (66)
Therefore, $\rho_{jj}=1$ as expected from the correlation coefficient between $X_{j}(1)$ and $X_{j}(1)$.
Now, let us deal with the more interesting case, i.e., $\rho_{12}$. We have the following
$\sigma_{1}\sigma_{2}\rho_{12}$ $=4b_{1}\left(\frac{H_{1}+H_{2}}{2}\right)\Re(c_{12})$
$=4\frac{\Gamma(2-(H_{1}+H_{2}))\cos\left(\frac{H_{1}+H_{2}}{2}\pi\right)}{(H_{1}+H_{2})(1-(H_{1}+H_{2}))}\frac{1}{2\pi}\Gamma\left(H_{1}+\frac{1}{2}\right)\Gamma\left(H_{2}+\frac{1}{2}\right)\rho\sigma_{1}\sigma_{2}a_{H_{1}}a_{H_{2}}\cos\left(\frac{(H_{2}-H_{1})\pi}{2}\right)$
$=4\rho\sigma_{1}\sigma_{2}\frac{\overbrace{\Gamma(1-(H_{1}+H_{2}))}^{\Gamma(z)\Gamma(1-z)=\frac{\pi}{\sin(\pi z)}}\cos\left(\frac{H_{1}+H_{2}}{2}\pi\right)}{H_{1}+H_{2}}\frac{1}{2\pi}\Gamma\left(H_{1}+\frac{1}{2}\right)\Gamma\left(H_{2}+\frac{1}{2}\right)a_{H_{1}}a_{H_{2}}\cos\left(\frac{(H_{2}-H_{1})\pi}{2}\right)$
$=2\rho\sigma_{1}\sigma_{2}\frac{\pi}{\sin((H_{1}+H_{2})\pi)\Gamma(H_{1}+H_{2})}\frac{\cos\left(\frac{H_{1}+H_{2}}{2}\pi\right)}{H_{1}+H_{2}}\frac{1}{\pi}\Gamma\left(H_{1}+\frac{1}{2}\right)\Gamma\left(H_{2}+\frac{1}{2}\right)\underbrace{a_{H_{1}}}_{=\frac{\sqrt{\Gamma(2H_{1}+1)\sin(H_{1}\pi)}}{\Gamma(H_{1}+\frac{1}{2})}}a_{H_{2}}\cos\left(\frac{(H_{2}-H_{1})\pi}{2}\right)$
$=\rho\sigma_{1}\sigma_{2}\frac{\sqrt{\Gamma(2H_{1}+1)\Gamma(2H_{2}+1)\sin(H_{1}\pi)\sin(H_{2}\pi)}}{\Gamma(H_{1}+H_{2}+1)\sin\left(\frac{H_{1}+H_{2}}{2}\pi\right)}\cos\left(\frac{(H_{2}-H_{1})\pi}{2}\right).$ (67)
Thus, the cross-correlation $\rho_{12}$ is given by
$\rho_{12}=\rho\frac{\sqrt{\Gamma(2H_{1}+1)\Gamma(2H_{2}+1)\sin(H_{1}\pi)\sin(H_{2}\pi)}}{\Gamma(H_{1}+H_{2}+1)\sin\left(\frac{H_{1}+H_{2}}{2}\pi\right)}\cos\left(\frac{(H_{2}-H_{1})\pi}{2}\right).$ (68)
Note that, for the special case $H_{1}=H_{2}$, the cross-correlation $\rho_{12}$ is equal to the correlation between noises, i.e., $\rho_{12}=\rho$. Lastly, to have the full covariance structure, we need to calculate coefficients $\eta_{jk}$. First, as $\Im(c_{jj})=0$ ($C$ being a Hermitian matrix), we have $\eta_{jj}=0\text{ for }j=1,2$. The more interesting case is $j\neq k$, which after similar calculations as for $\rho_{12}$ leads to
$\eta_{12}=\rho\frac{\sqrt{\Gamma(2H_{1}+1)\Gamma(2H_{2}+1)\sin(H_{1}\pi)\sin(H_{2}\pi)}}{\Gamma(H_{1}+H_{2}+1)\cos\left(\frac{H_{1}+H_{2}}{2}\pi\right)}\sin\left(\frac{(H_{2}-H_{1})\pi}{2}\right).$ (69)
It can be expressed as
$\eta_{12}=\rho_{12}\tan\left(\pi\frac{H_{2}-H_{1}}{2}\right)\tan\left(\pi\frac{H_{1}+H_{2}}{2}\right)=-\rho_{12}\tan\left(\pi\frac{H_{1}-H_{2}}{2}\right)\tan\left(\pi\frac{H_{1}+H_{2}}{2}\right),$ (70)
which coincides with the results for the causal process given in Section 3.3 of [1].
Appendix B Construction of a well-balanced 2D fBm
First, let us consider a one-dimensional well-balanced fractional Brownian motion given in Definition 3 below.
Defintition 3 (Well-balanced 1D fBm).
Well-balanced fractional Brownian motion $X^{*}(t),t\geq 0,$ is given by the following time-domain representation
$X^{*}(t)=a^{*}_{H}\int_{\mathbb{R}}((t-u)_{+}^{H-\frac{1}{2}}-(-u)_{+}^{H-\frac{1}{2}})+((t-u)_{-}^{H-\frac{1}{2}}-(-u)_{-}^{H-\frac{1}{2}}))B(\mathrm{d}u),$ (71)
where $H\in(0,1)$ and $a^{*}_{H}>0$.
Theorem 6.
The constant $a^{*}_{H}$ for which $\langle X^{*2}(t)\rangle=1$ for the well-balanced 1D fBm is given by
$a^{*2}_{H}$ $=\frac{2H(1-2H)\pi}{8\Gamma(2-2H)\cos(H\pi)\Gamma^{2}(H+\frac{1}{2})\cos^{2}\left(\frac{\pi(H-\frac{1}{2})}{2}\right)}.$
Proof.
Let us recall the definition of the function given in Eq. (1)
$f_{\pm}(x;t,\beta)=(t-x)_{\pm}^{\beta}-(-x)_{\pm}^{\beta}.$
In order to calculate the autocovariance of the process, let us first utilize Itô’s lemma. We have
$\langle X^{*}(t)X^{*}(s)\rangle$ $=a^{*2}_{H}\int_{\mathbb{R}}(f_{+}(u;t,D)+f_{-}(u;t,D))(f_{+}(u;s,D)+f_{-}(u;s,D))\mathrm{d}u$
$\stackrel{{\scriptstyle\textrm{Plancherel}}}{{=}}\frac{a^{*2}_{H}}{2\pi}\int_{\mathbb{R}}(\widehat{f_{+}}(\xi;t,D)+\widehat{f_{-}}(\xi;t,D))\overline{(\widehat{f_{+}}(\xi;s,D)+\widehat{f_{-}}(\xi;s,D))}\mathrm{d}\xi,$ (72)
where $\overline{z}$ is a complex conjugate of $z$ and $\widehat{f_{\pm}}(\xi;t,D)$ is a Fourier transform of $f_{\pm}$. Thus, it is given by [16] (formulae 3.76.4 and 3.76.9)
$\widehat{f_{\pm}}(\xi;t,D)\equiv\int_{\mathbb{R}}e^{iu\xi}f(u;t,D)\mathrm{d}u=\frac{e^{itx}-1}{ix|x|^{D}}\Gamma(D+1)e^{\mp i\mathrm{sign}(x)\frac{\pi D}{2}}$ (73)
and
$\widehat{f_{+}}(\xi;t,D)+\widehat{f_{-}}(\xi;t,D)=\frac{e^{itx}-1}{ix|x|^{D}}\Gamma(D+1)2\cos^{2}\left(\frac{\pi D}{2}\right).$ (74)
Consequently, the autocovariance function becomes
$\langle X^{*}(t)X^{*}(s)\rangle$ $=\frac{a^{*2}_{H}}{2\pi}\int_{\mathbb{R}}\frac{e^{itx}-1}{ix|x|^{D}}\frac{e^{-isx}-1}{-ix|x|^{D}}\Gamma^{2}(D+1)4\cos^{2}\frac{\pi D}{2}\mathrm{d}x=$
$=4\frac{a^{*2}_{H}}{2\pi}\Gamma^{2}(D+1)\cos^{2}\left(\frac{\pi D}{2}\right)\int_{\mathbb{R}}\frac{(e^{itx}-1)(e^{-isx}-1)}{|x|^{2(D+1)}}\mathrm{d}x.$ (75)
When $t=s=1$ we have
$\langle X^{*2}(1)\rangle$ $=4\frac{a^{*2}_{H}}{2\pi}\Gamma^{2}(D+1)\cos^{2}\left(\frac{\pi D}{2}\right)\int_{\mathbb{R}}\frac{(e^{ix}-1)(e^{-ix}-1)}{|x|^{2(D+1)}}\mathrm{d}x$
$=\frac{4a^{*2}_{H}}{\pi}\Gamma^{2}(D+1)\cos^{2}\left(\frac{\pi D}{2}\right)\int_{\mathbb{R}}\frac{1-\cos x}{|x|^{2(D+1)}}\mathrm{d}x$
$=\frac{a^{*2}_{H}}{\pi}\Gamma^{2}(D+1)\cos^{2}\left(\frac{\pi D}{2}\right)\frac{4\pi}{\Gamma(2(D+1))\sin((D+\frac{1}{2})\pi)}$
$=\frac{a^{*2}_{H}}{\pi}\Gamma^{2}(H+\frac{1}{2})\cos^{2}\left(\frac{\pi(H-\frac{1}{2})}{2}\right)\frac{4\pi}{\Gamma(2H+1)\sin(H\pi)}.$ (76)
To ensure that the variance of the process $X^{*}(t)$ is equal to 1 at time $t=1$, the constant $a_{H}$ has to be as follows
$a^{*2}_{H}$ $=\frac{2H\Gamma(2H)\sin(\pi H)}{4\Gamma^{2}(H+\frac{1}{2})\cos^{2}\left(\frac{\pi(H-\frac{1}{2})}{2}\right)}$
$=\frac{2H(1-2H)\pi}{8\Gamma(2-2H)\cos(H\pi)\Gamma^{2}(H+\frac{1}{2})\cos^{2}\left(\frac{\pi(H-\frac{1}{2})}{2}\right)}.$ (77)
Appendix C Proof of Theorem 1, well-balanced case and remarks
Proof.
Again, looking at the general form given in Eq. (51), we can calculate matrices $M_{+},M_{-}$ and then, $A_{1},A_{2}$ and $C$. It is clear that matrices $M_{+}$ and $M_{-}$ are the same, and given by
$M_{\pm}=\begin{bmatrix}a^{*}_{H_{1}}\sigma_{1}&0\\ a^{*}_{H_{2}}\rho\sigma_{2}&a^{*}_{H_{2}}\sqrt{1-\rho^{2}}\sigma_{2}\end{bmatrix}$. (78)
Thus, by the relations given in Eqs. (54)-(55) we have
$A_{1}$ $=0\cdot I,$ (79)
$A_{2}$ $=\sqrt{\frac{2}{\pi}}\cos\frac{D\pi}{2}\Gamma(D+I)M_{+}=\begin{bmatrix}a^{*}_{H_{1}}\sigma_{1}\cos\left(\frac{(H_{1}-\frac{1}{2})\pi}{2}\right)\Gamma\left(H_{1}+\frac{1}{2}\right)&0\\ a^{*}_{H_{2}}\rho\sigma_{2}\cos\left(\frac{(H_{2}-\frac{1}{2})\pi}{2}\right)\Gamma\left(H_{2}+\frac{1}{2}\right)&a^{*}_{H_{2}}\sqrt{1-\rho^{2}}\sigma_{2}\cos\left(\frac{(H_{2}-\frac{1}{2})\pi}{2}\right)\Gamma\left(H_{2}+\frac{1}{2}\right)\end{bmatrix}$. (80)
In the case of well-balanced 2D fBm, $A=A_{1}+iA_{2}$ can be written as
$A$ $=iA_{2}=i\begin{bmatrix}a^{*}_{H_{1}}\sigma_{1}\cos\left(\frac{(H_{1}-\frac{1}{2})\pi}{2}\right)\Gamma\left(H_{1}+\frac{1}{2}\right)&0\\ a^{*}_{H_{2}}\rho\sigma_{2}\cos\left(\frac{(H_{2}-\frac{1}{2})\pi}{2}\right)\Gamma\left(H_{2}+\frac{1}{2}\right)&a^{*}_{H_{2}}\sqrt{1-\rho^{2}}\sigma_{2}\cos\left(\frac{(H_{2}-\frac{1}{2})\pi}{2}\right)\Gamma\left(H_{2}+\frac{1}{2}\right)\end{bmatrix}$. (81)
Since $A=A_{1}+iA_{2}=iA_{2}$ and $C=AA^{*}$, elements of $C$ are given by
$\begin{cases}c_{jj}&=\frac{2}{\pi}\sigma^{2}_{j}\cos^{2}\left(\frac{(H_{j}-0.5)\pi}{2}\right)\Gamma^{2}(H_{j}+0.5)a^{*2}_{H_{j}},\\ c_{jk}&=\frac{2}{\pi}\rho\sigma_{j}\sigma_{k}\cos\left(\frac{(H_{j}-0.5)\pi}{2}\right)\cos\left(\frac{(H_{k}-0.5)\pi}{2}\right)\Gamma(H_{j}+0.5)\Gamma(H_{k}+0.5)a^{*}_{H_{j}}a^{*}_{H_{k}},\quad j\neq k,\end{cases}$ (82)
for $j,k=1,2$. We note that $C$ is a real matrix, in contrast to its equivalent in the causal 2D fBm. The consequence of that is that $\eta_{jk}=0$ for all $j,k=1,2$, and therefore this is a time-reversible version of a 2D fBm. Thus, the covariance structure of two-dimensional well-balanced 2D fBm simplifies to
$\langle X^{*}_{j}(t)X^{*}_{k}(s)\rangle=\frac{\sigma_{j}\sigma_{k}\rho_{jk}}{2}\left(|t|^{H_{j}+H_{k}}+|s|^{H_{j}+H_{k}}-|t-s|^{H_{j}+H_{k}}\right).$ (83)
We can also calculate coefficients $\rho_{jk}$ present in the covariance structure
$\sigma_{j}^{2}\rho_{jj}$ $=4b_{1}(H_{j})\Re(c_{jj})=4\underbrace{\frac{\Gamma(2-2H_{j})\cos(H_{j}\pi)}{2H_{j}(1-2H_{j})}}_{b_{1}(H_{j})}\underbrace{\frac{2}{\pi}\sigma_{j}^{2}a_{H_{j}}^{2}\Gamma^{2}\left(H_{j}+\frac{1}{2}\right)\cos^{2}\left(\frac{(H_{j}-\frac{1}{2})\pi}{2}\right)}_{c_{jj}}$
$=4\frac{\Gamma(2-2H_{j})\cos(H_{j}\pi)}{2H_{j}(1-2H_{j})}\frac{2}{\pi}\sigma_{j}^{2}\Gamma\left(H_{j}+\frac{1}{2}\right)\cos^{2}\left(\frac{(H_{j}-\frac{1}{2})\pi}{2}\right)\frac{2H_{j}(1-2H_{j})\pi}{8\Gamma(2-2H_{j})\cos(\pi H_{j})\Gamma^{2}\left(H_{j}+\frac{1}{2}\right)\cos^{2}\left(\frac{(H_{j}-\frac{1}{2})\pi}{2}\right)}=\sigma_{j}^{2}.$ (84)
As expected $\rho_{jj}=1$, since it is a correlation of $X^{*}_{j}(1)$ with itself. For more interesting case, i.e., for $\rho_{12}$, we have
$\sigma_{1}\sigma_{2}\rho_{12}$ $=4b_{1}\left(\frac{H_{1}+H_{2}}{2}\right)\Re(c_{12})=$
$=4\underbrace{\frac{\Gamma(2-(H_{1}+H_{2}))\cos\frac{(H_{1}+H_{2})\pi}{2}}{(H_{1}+H_{2})(1-(H_{1}+H_{2}))}}_{b_{1}\left(\frac{H_{1}+H_{2}}{2}\right)}\underbrace{\frac{2}{\pi}\cos\left(\frac{(H_{1}-\frac{1}{2})\pi}{2}\right)\Gamma\left(H_{1}+\frac{1}{2}\right)\cos\left(\frac{(H_{2}-\frac{1}{2})\pi}{2}\right)\Gamma\left(H_{2}+\frac{1}{2}\right)\sigma_{1}\sigma_{2}a^{*}_{H_{1}}a^{*}_{H_{2}}\rho}_{\Re(c_{12})}$
$=\frac{8}{\pi}\sigma_{1}\sigma_{2}\rho\frac{\Gamma(2-(H_{1}+H_{2}))\cos\frac{(H_{1}+H_{2})\pi}{2}}{(H_{1}+H_{2})(1-(H_{1}+H_{2}))}\frac{\cos\left(\frac{(H_{1}-\frac{1}{2})\pi}{2}\right)\Gamma\left(H_{1}+\frac{1}{2}\right)\cos\left(\frac{(H_{2}-\frac{1}{2})\pi}{2}\right)\Gamma\left(H_{2}+\frac{1}{2}\right)\sqrt{2H_{1}(1-2H_{1})}\sqrt{2H_{2}(1-2H_{2})}\pi}{8\sqrt{\Gamma(2-2H_{1})\cos(\pi H_{1})}\sqrt{\Gamma(2-2H_{2})\cos(\pi H_{2})}\Gamma\left(H_{1}+\frac{1}{2}\right)\cos\left(\frac{(H_{1}-\frac{1}{2})\pi}{2}\right)\Gamma\left(H_{2}+\frac{1}{2}\right)\cos\left(\frac{(H_{2}-\frac{1}{2})\pi}{2}\right)}$
$=\sigma_{1}\sigma_{2}\rho\frac{\cos\left(\frac{(H_{1}+H_{2})\pi}{2}\right)}{\Gamma(H_{1}+H_{2}+1)\sin((H_{1}+H_{2})\pi)}\frac{\sqrt{\Gamma(2H_{1}+1)\Gamma(2H_{2}+1)\sin(2H_{1}\pi)\sin(2H_{2}\pi)}}{\sqrt{\cos(H_{1}\pi)\cos(H_{2}\pi)}}$
$=\sigma_{1}\sigma_{2}\rho\frac{\sqrt{\Gamma(2H_{1}+1)\Gamma(2H_{2}+1)}}{\Gamma(H_{1}+H_{2}+1)}\frac{\sqrt{\sin(H_{1}\pi)\sin(H_{2}\pi)}}{\sin\left(\frac{(H_{1}+H_{2})\pi}{2}\right)}.$ (85)
And thus,
$\rho_{12}=\rho\frac{\sqrt{\Gamma(2H_{1}+1)\Gamma(2H_{2}+1)}}{\Gamma(H_{1}+H_{2}+1)}\frac{\sqrt{\sin(H_{1}\pi)\sin(H_{2}\pi)}}{\sin\left(\frac{(H_{1}+H_{2})\pi}{2}\right)},$ (86)
which differs from the one given in the causal case missing a $\cos\left(\frac{(H_{2}-H_{1})\pi}{2}\right)$ factor. Similarly, the condition that
$\rho_{12}^{2}\leq\left(\frac{\sqrt{\Gamma(2H_{1}+1)\Gamma(2H_{2}+1)}}{\Gamma(H_{1}+H_{2}+1)}\frac{\sqrt{\sin(H_{1}\pi)\sin(H_{2}\pi)}}{\sin\left(\frac{(H_{1}+H_{2})\pi}{2}\right)}\right)^{2}$ (87)
is the one present in the literature [42] that guarantees the proper covariance structure.
Appendix D Proof of Theorem 3, spectral content of the increment process of 2D fBm
Proof.
First, following [42], we know the spectral representation of the increments
$\Delta\mathbf{Z}(n)\equiv\mathbf{Z}(n+1)-\mathbf{Z}(n)=\int_{\mathbb{R}}\frac{e^{i(n+1)x}-e^{inx}}{ix}\left(x_{+}^{-D}A+x_{-}^{-D}\overline{A}\right)\widehat{B}(\mathrm{d}x),$ (88)
where $A$ is given in Eq. (58), $D=H-\frac{1}{2}I$ and $\tilde{B}$ is Gaussian spectral measure with $\left\langle|\widehat{B}(\mathrm{d}x)|^{2}\right\rangle=\mathrm{d}x$ control measure. Now we can write
$\left\langle\Delta\mathbf{Z}(k)\Delta\mathbf{Z}(0)\right\rangle$ $=\left\langle\int_{\mathbb{R}}\int_{\mathbb{R}}\frac{e^{i(k+1)x}-e^{ikx}}{ix}\overline{\left(\frac{e^{iy}-1}{iy}\right)}\left(x_{+}^{-D}A+x_{-}^{-D}\overline{A}\right)\cdot\left(y_{+}^{-D}A+y_{-}^{-D}\overline{A}\right)^{*}\widehat{B}(\mathrm{d}x)\widehat{B}^{*}(\mathrm{d}y)\right\rangle$
$=\int_{\mathbb{R}}\frac{e^{i(k+1)x}-e^{ikx}}{ix}\overline{\left(\frac{e^{ix}-1}{ix}\right)}\left(x_{+}^{-D}A+x_{-}^{-D}\overline{A}\right)\cdot\left(x_{+}^{-D}A+x_{-}^{-D}\overline{A}\right)^{*}\mathrm{d}x$
$=\sum_{n=-\infty}^{\infty}\int_{-\pi+2n\pi}^{\pi+2n\pi}\frac{e^{i(k+1)x}-e^{ikx}}{ix}\overline{\frac{e^{ix}-1}{ix}}\left(x_{+}^{-D}A+x_{-}^{-D}\overline{A}\right)\cdot\left(x_{+}^{-D}A+x_{-}^{-D}\overline{A}\right)^{*}\mathrm{d}x$
[…]
$=\int_{-\pi}^{\pi}\sum_{n=-\infty}^{\infty}e^{ik(f+2n\pi)}\frac{|e^{i(f+2n\pi)}-1|^{2}}{(f+2n\pi)^{2}}\left((f+2n\pi)_{+}^{-D}A+(f+2n\pi)_{-}^{-D}\overline{A}\right)\cdot\left((f+2n\pi)_{+}^{-D}A+(f+2n\pi)_{-}^{-D}\overline{A}\right)^{*}\mathrm{d}f$
$=\int_{-\pi}^{\pi}e^{ikf}|e^{if}-1|^{2}\sum_{n=-\infty}^{\infty}\left((f+2n\pi)_{+}^{-D}AA^{*}(f+2n\pi)_{+}^{-D}+(f+2n\pi)_{-}^{-D}\overline{A}\overline{A^{*}}(f+2n\pi)_{-}\right)/(f+2n\pi)^{2}\mathrm{d}f.$
Remembering that $AA^{*}=C$ (elements of $C$ are given in (59) and (82)), and taking Remark 1.3.6 and 1.3.8 [42] we obtain the final result for the power spectral density of $\Delta\mathbf{Z}$.
Considering element-wise power spectral density, we have
$S_{\Delta\mathbf{Z},jk}(f)$ $=|e^{if}-1|^{2}\sum_{n=-\infty}^{\infty}\left[(f+2n\pi)_{+}^{-(d_{j}+d_{k})}c_{jk}+(f+2n\pi)_{-}^{-(d_{j}+d_{k})}\overline{c_{jk}}\right]/(f+2n\pi)^{2}$
$=|e^{if}-1|^{2}\sum_{n=-\infty}^{\infty}\left[(f+2n\pi)_{+}^{-(H_{j}+H_{k}+1)}c_{jk}+(f+2n\pi)_{-}^{-(H_{j}+H_{k}+1)}\overline{c_{jk}}\right]/(f+2n\pi)^{2},$
as $d_{j}=H_{j}-1/2$.
Moreover, by simple calculations we have the asymptotics as $f\to 0$
$S_{\Delta\mathbf{Z},jk}(f)\sim c_{jk}f^{(-H_{j}-H_{k}+1)}.$
Appendix E Proof of Theorem 4, spectral content of 2D fBm trajectory
Proof.
Power spectral density for the trajectory $\mathbf{Z}(t),0\leq t\leq T$, is given by
$S_{jk}(f,T)\equiv\frac{1}{T}\int_{0}^{T}e^{ift}Z_{j}(t)\,\mathrm{d}t\overline{\int_{0}^{T}e^{ifs}Z_{k}(s)\,\mathrm{d}s}.$ (89)
It is a random variable depending both on the considered frequency $f$, time horizon $T$, and the trajectory $\mathbf{Z}(t)$. Since we want to consider the ensemble-averaged power spectral density, we have
$\langle S_{jk}(f,T)\rangle$ $=\frac{1}{T}\left\langle\int_{0}^{T}e^{ift}Z_{j}(t)\,\mathrm{d}t\overline{\int_{0}^{T}e^{ifs}Z_{k}(s)\,\mathrm{d}s}\right\rangle=\ |\textrm{by Fubini theorem}|$
$=\frac{1}{T}\int_{0}^{T}\int_{0}^{T}e^{if(t-s)}\left\langle Z_{j}(t)Z_{k}(s)\right\rangle\,\mathrm{d}t\,\mathrm{d}s$
$=\frac{1}{T}\int_{0}^{T}\int_{0}^{T}e^{if(t-s)}\gamma_{jk}(t,s)\,\mathrm{d}t\,\mathrm{d}s,$
where $\gamma_{jk}(t,s)$ is the covariance function for $Z_{j}$ and $Z_{k}$. For simplicity, we write the general form of this function with weighting function $w_{jk}$ that has explicit form for both causal and well-balanced cases (Eq. (13)).
Here, we consider the case when $H_{j}+H_{k}\neq 1$ or the considered model is well-balanced 2D fBm (so that, there is no $\log$ part in the cross covariance function). By rescaling the integrals’ variables $\frac{t}{T}\rightarrow x,\frac{s}{T}\rightarrow y$ and by introducing $\tilde{\omega}=fT$, we have
$\langle S_{jk}(f,T)\rangle$ $=T^{H_{j}+H_{k}+1}\frac{\sigma_{j}\sigma_{k}}{2}\int_{0}^{1}\int_{0}^{1}e^{i\tilde{\omega}(x-y)}\left(w_{jk}(x)x^{H_{j}+H_{k}}+w_{jk}(-y)y^{H_{j}+H_{k}}-w_{jk}(x-y)|x-y|^{H_{j}+H_{k}}\right)\,\mathrm{d}x\,\mathrm{d}y$ (90)
We calculate each element separately. Let us introduce the notation
$\mathbb{I}$ $\equiv\int_{0}^{1}\int_{0}^{1}e^{i\tilde{\omega}(x-y)}w_{jk}(x)x^{H_{j}+H_{k}}\mathrm{d}y\,\mathrm{d}x,$ (91)
$\mathbb{II}$ $\equiv\int_{0}^{1}\int_{0}^{1}e^{i\tilde{\omega}(x-y)}w_{jk}(-y)y^{H_{j}+H_{k}}\mathrm{d}y\,\mathrm{d}x,$ (92)
$\mathbb{III}$ $\equiv\int_{0}^{1}\int_{0}^{1}e^{i\tilde{\omega}(x-y)}w_{jk}(x-y)|x-y|^{H_{j}+H_{k}}\mathrm{d}y\,\mathrm{d}x.$ (93)
Then the resulting integral is equal to
$\langle S_{jk}(f,T)\rangle$ $=T^{H_{j}+H_{k}+1}\frac{\sigma_{j}\sigma_{k}}{2}(\mathbb{I}+\mathbb{II}-\mathbb{III}).$ (94)
Let us start with $\mathbb{I}$
$\mathbb{I}$ $\equiv\int_{0}^{1}\int_{0}^{1}e^{i\tilde{\omega}(x-y)}w_{jk}(x)x^{H_{j}+H_{k}}\mathrm{d}y\,\mathrm{d}x$
$=\int_{0}^{1}e^{i\tilde{\omega}x}\frac{1-e^{-i\tilde{\omega}}}{i\tilde{\omega}}x^{H_{j}+H_{k}}\underbrace{w_{jk}(x)}_{\text{constant }\forall x>0}\,\mathrm{d}x$
$=w_{jk}(1)\frac{1-e^{-i\tilde{\omega}}}{i\tilde{\omega}}\left[\int_{0}^{1}\cos(\tilde{\omega}x)+i\sin(\tilde{\omega}x))x^{H_{j}+H_{k}}\,\mathrm{d}x\right].$
Now, let us consider $\mathbb{II}$
$\mathbb{II}$ $\equiv\int_{0}^{1}\int_{0}^{1}e^{i\tilde{\omega}(x-y)}w_{jk}(-y)y^{H_{j}+H_{k}}\mathrm{d}x\,\mathrm{d}y$
$=\int_{0}^{1}e^{-i\tilde{\omega}y}\frac{e^{i\tilde{\omega}}-1}{i\tilde{\omega}}y^{H_{j}+H_{k}}\underbrace{w_{jk}(-y)}_{\text{constant }\forall y>0}\,\mathrm{d}y$
$=w_{jk}(-1)\frac{e^{i\tilde{\omega}}-1}{i\tilde{\omega}}\left[\int_{0}^{1}\cos(\tilde{\omega}x)-i\sin(\tilde{\omega}x))y^{H_{j}+H_{k}}\,\mathrm{d}x\right].$
Summing up the first two parts, we have
$\mathbb{I}+\mathbb{II}=$ $\int_{0}^{1}\cos(\tilde{\omega}x)x^{H_{j}+H_{k}}\,\mathrm{d}x\left[w_{jk}(1)\frac{1-e^{-i\tilde{\omega}}}{i\tilde{\omega}}-w_{jk}(-1)\frac{1-e^{i\tilde{\omega}}}{i\tilde{\omega}}\right$] (95)
$+\int_{0}^{1}\sin(\tilde{\omega}y)y^{H_{j}+H_{k}}\,\mathrm{d}y\left[w_{jk}(1)\frac{1-e^{-i\tilde{\omega}}}{i\tilde{\omega}}+w_{jk}(-1)\frac{1-e^{i\tilde{\omega}}}{i\tilde{\omega}}\right$]. (96)
Reintroducing functions $\mathcal{C}_{jk}$ and $\mathcal{S}_{jk}$ as
[…]
and utilizing that $w_{jk}(x)=\rho_{jk}-\mathrm{sign}(x)\eta_{jk}$ we can simplify the expression
$\mathbb{I}+\mathbb{II}=$ $\mathcal{C}_{jk}\left[(\rho_{jk}-\eta_{jk})\frac{1-\cos(\tilde{\omega})+i\sin(\tilde{\omega})}{i\tilde{\omega}}-(\rho_{jk}+\eta_{jk})\frac{1-\cos(\tilde{\omega})-i\sin(\tilde{\omega})}{i\tilde{\omega}}\right$]
$+\mathcal{S}_{jk}\left[(\rho_{jk}-\eta_{jk})\frac{1-\cos(\tilde{\omega})+i\sin(\tilde{\omega})}{i\tilde{\omega}}-(\rho_{jk}+\eta_{jk})\frac{1-\cos(\tilde{\omega})-i\sin(\tilde{\omega})}{i\tilde{\omega}}\right$]
$=$ $2\rho_{jk}\left(\frac{\sin\tilde{\omega}}{\tilde{\omega}}\mathcal{C}_{jk}+\frac{1-\cos\tilde{\omega}}{\tilde{\omega}}\mathcal{S}_{jk}\right)+2i\eta_{jk}\left(\frac{1-\cos\tilde{\omega}}{\tilde{\omega}}\mathcal{C}_{jk}-\frac{\sin\tilde{\omega}}{\tilde{\omega}}\mathcal{S}_{jk}\right)$ (99)
For the third part, we have
$\mathbb{III}$ $\equiv\int_{0}^{1}\int_{0}^{1}e^{i\tilde{\omega}(x-y)}w_{jk}(x-y)|x-y|^{H_{j}+H_{k}}\mathrm{d}x\,\mathrm{d}y=\quad|\textrm{substitute }u=x-y,\ z=x|$
$=\int_{-1}^{0}du\int_{0}^{u+1}dzw_{jk}(u)e^{i\tilde{\omega}u}|u|^{H_{j}+H_{k}}+\int_{0}^{1}du\int_{u}^{1}dzw_{jk}(u)e^{i\tilde{\omega}u}u^{H_{j}+H_{k}}$
$=\int_{-1}^{0}duw_{jk}(u)e^{i\tilde{\omega}u}|u|^{H_{j}+H_{k}}(1+u)+\int_{0}^{1}duw_{jk}(u)e^{i\tilde{\omega}u}u^{H_{j}+H_{k}}(1-u)$
$=\int_{0}^{1}du\underbrace{w_{jk}(-u)}_{=\rho_{jk}+\eta_{jk}}e^{-i\tilde{\omega}u}u^{H_{j}+H_{k}}(1-u)+\int_{0}^{1}du\underbrace{w_{jk}(u)}_{=\rho_{jk}-\eta_{jk}}e^{i\tilde{\omega}u}u^{H_{j}+H_{k}}(1-u)$
$=\rho_{jk}\int_{0}^{1}\underbrace{\left(e^{-i\tilde{\omega}u}+e^{i\tilde{\omega}u}\right)}_{=2\cos(\tilde{\omega}u)}u^{H_{j}+H_{k}}(1-u)du-\eta_{jk}\int_{0}^{1}\underbrace{\left(e^{i\tilde{\omega}u}-e^{-i\tilde{\omega}u}\right)}_{=2i\sin(\tilde{\omega}u)}u^{H_{j}+H_{k}}(1-u)du$
$=2\rho_{jk}\left(\mathcal{C}_{jk}-\int_{0}^{1}\cos(\tilde{\omega}u)u^{H_{j}+H_{k}+1}du\right)-2i\eta_{jk}\left(\mathcal{S}_{jk}-\int_{0}^{1}\sin(\tilde{\omega}u)u^{H_{j}+H_{k}+1}du\right).$
We see that
$\int_{0}^{1}\cos(\tilde{\omega}u)u^{H_{j}+H_{k}+1}du$ $=\frac{\mathrm{d}}{\mathrm{d}\tilde{\omega}}\int_{0}^{1}\sin(\tilde{\omega}u)u^{H_{j}+H_{k}}du=\frac{\mathrm{d}}{\mathrm{d}\tilde{\omega}}\mathcal{S}_{jk},$ (100)
$\int_{0}^{1}\sin(\tilde{\omega}u)u^{H_{j}+H_{k}+1}du$ $=-\frac{\mathrm{d}}{\mathrm{d}\tilde{\omega}}\int_{0}^{1}\cos(\tilde{\omega}u)u^{H_{j}+H_{k}}du=-\frac{\mathrm{d}}{\mathrm{d}\tilde{\omega}}\mathcal{C}_{jk},$ (101)
or, via integration by parts we obtain
$\int_{0}^{1}\cos(\tilde{\omega}u)u^{H_{j}+H_{k}+1}du=\frac{\sin\tilde{\omega}}{\tilde{\omega}}-\frac{H_{j}+H_{k}+1}{\tilde{\omega}}\mathcal{S}_{jk},$ (102)
$\int_{0}^{1}\sin(\tilde{\omega}u)u^{H_{j}+H_{k}+1}du=\frac{-\cos\tilde{\omega}}{\tilde{\omega}}+\frac{H_{j}+H_{k}+1}{\tilde{\omega}}\mathcal{C}_{jk}.$ (103)
We then arrive at the final expression
$\mathbb{I}+\mathbb{II}-\mathbb{III}=$ $2\rho_{jk}\left(\left[\frac{\sin\tilde{\omega}}{\tilde{\omega}}-1\right]\mathcal{C}_{jk}+\frac{1-\cos\tilde{\omega}}{\tilde{\omega}}\mathcal{S}_{jk}+\frac{\mathrm{d}}{\mathrm{d}\tilde{\omega}}\mathcal{S}_{jk}\right)$
$+2i\eta_{jk}\left(\frac{1-\cos\tilde{\omega}}{\tilde{\omega}}\mathcal{C}_{jk}-\left[\frac{\sin\tilde{\omega}}{\tilde{\omega}}+1\right]\mathcal{S}_{jk}-\frac{\mathrm{d}}{\mathrm{d}\tilde{\omega}}\mathcal{C}_{jk}\right),$ (104)
or equivalently
[…]
$+2i\eta_{jk}\left(\frac{1-\cos\tilde{\omega}+H_{j}+H_{k}+1}{\tilde{\omega}}\mathcal{C}_{jk}-\left[\frac{\sin\tilde{\omega}}{\tilde{\omega}}+1\right]\mathcal{S}_{jk}\right).$ (105)
Appendix F Proof of Theorem 5, asymptotics of ensemble-averaged PSD
Proof.
Per [25] we have the asymptotic behaviour of functions $\mathcal{C}_{jk}$ and $\mathcal{S}_{jk}$
[…]
Thus, grouping the elements by their powers, we have (as $\tilde{\omega}\to 0$)
$\Re\langle S_{jk}(\tilde{\omega},T)\rangle$ $\simeq T^{H_{j}+H_{k}+1}\sigma_{j}\sigma_{k}\rho_{jk}\left\{\left[\frac{1}{\tilde{\omega}^{2}}-\frac{(H_{j}+H_{k})\sin\tilde{\omega}}{\tilde{\omega}^{3}}+O\left(\frac{1}{\tilde{\omega}^{4}}\right)\right]\right$. (108)
[…]
$\Im\langle S_{jk}(\tilde{\omega},T)\rangle$ $\simeq T^{H_{j}+H_{k}+1}\sigma_{j}\sigma_{k}\eta_{jk}\left\{\left[\frac{2}{\tilde{\omega}^{2}}-\frac{(H_{j}+H_{k})\cos\tilde{\omega}(3-\cos\tilde{\omega})}{\tilde{\omega}^{3}}+O\left(\frac{1}{\tilde{\omega}^{4}}\right)\right]\right$.
[…]